0.05x+0.25y=44
0.15x+0.05y=48
Elimination method:

Answer by sowmya(32)

0.05x+0.25y=44
0.15x+0.05y=48
Elimination method:
eq 1:
5x +25y = 4400
x +5y = 880
x = 880-5y
eq 2:
15x+5y = 4800
3x+y = 960
3(880-5y) +y =960
2640 -15y + y = 960
14y = 2580
y = 2580/14 = 1290/7
x = (5860-6450)/7 = - 590/7